Question: A review of front-desk records at Surfside Hotel during the peak period reveals the following no-show data: 0 (with probability 0.10), 1 (with probability 0.30), 2 (with probability 0.40), 3 (with probability 0.20). The room rate per night is $75. In case of overbooking, the guest will be walked to a nearby hotel and Surfside will pay the room rate of $125 to accommodate the guest. According to the information, what is the overbooking cost Co and what is the critical fractile?

Question options:
a) 125 and 0.625
b) 75 and 0.375
c) 125 and 0.375
d) 50 and 0.6
e) 50 and 0.4
Answer #1

Note: Please rate the answer (upvote) and do ask if any doubt is there.

Answer:

  • Option : 50 and 0.6

Explanation:

  • Cost of understock = Cu = price of a room = $75
  • Cost of overstock = Co = cost of surf side- price = 125-75= $50 (Answer)
  • P = critical fractile = Cu / (Cu + Co) = 75/125= 0.60 (Answer)

Note: Please rate the answer (upvote) and do ask if any doubt is there.